OpenStudy (anonymous):

(x-2xlnx)/x^4 how can I find the x value?

OpenStudy (anonymous):

\[\frac{x-2x\ln(x)}{x^4}\] what are you trying to do?

OpenStudy (anonymous):

x−2xln(x)/x4=0 need to fond x value.

OpenStudy (anonymous):

ooooh set numerator = 0 and solve

OpenStudy (anonymous):

\[x=2x\ln(x)\] \[\frac{1}{2}=\ln(x)\] \[x=e^{\frac{1}{2}}\]

OpenStudy (anonymous):

thank you =)